# A new approach to principal–agent problems with volatility control

### v1

@@ 4 Illustrative Examples
In this section, we present two illustrative examples, in order to highlight the link between 2 and 4, in particular the one-to-one correspondence between the contract’s sensitivity parameter $\Gamma$ in 2 and the quadratic variation density process $\Sigma$ in 4. The introduction and resolution of these two examples is rather informal, we refer to the previous sections for a more rigorous formulation and the theoretical results.
@@ 4.1 A first-best case
We first consider a very simple example, in which the agent only controls the volatility of the output process. More precisely, we let
$X_{t}:=x_{0}+\int_{0}^{t}\nu_{s}\mathrm{d}W_{s},\;t\in[0,T],\;x_{0}\in\mathbb{R},$
where $W$ here is a one-dimensional Brownian Motion, and $\nu$ is an ‘admissible’ control chosen by the agent, namely satisfying appropriate integrability conditions and taking values in $[0,1]$. We further assume that, given a contract $\xi$, the agent wishes to maximise the following expected utility,
@@
We first recall that, since the principal observes $X$ in continuous time, she can also compute the associated quadratic variation process, namely $\langle X\rangle$. Here, such process follows the dynamics $\mathrm{d}\langle X\rangle_{t}=\nu_{t}^{2}\mathrm{d}t$, $t\in[0,T]$. In other words, the quadratic variation density process, denoted $\Sigma$ throughout the paper, satisfies $\Sigma_{t}=\nu_{t}^{2}$, $t\in[0,T]$, and since $\nu$ takes values in $[0,1]$, it is already clear that, by observing $\Sigma$ in continuous-time, the principal can deduce the agent’s effort $\nu_{t}=\sqrt{\Sigma_{t}}$, $t\in[0,T]$. This example therefore illustrates a first-best scenario, since from her observation, the principal can actually deduce the agent’s effort.
If one follows the approach developed by [9], the optimal form of contract is the following
@@
for $y_{0}\in\mathbb{R}$, $(Z,\Gamma)\in{\cal V}$, both taking values in $\mathbb{R}$, and where the agent’s Hamiltonian is defined by
${\cal H}_{\rm A}(\gamma):=\dfrac{1}{2}\sup_{u\in[0,1]}\big{\{}\gamma u^{2}-u^{-2}\big{\}},\;\gamma\in\mathbb{R}.$
@@
If we follow our approach, we first need to fix the quadratic variation density process $\Sigma\in{\cal S}$. Note here that, since $\Sigma_{t}=|\nu_{t}|^{2}$, $t\in[0,T]$, and $\nu$ takes values in $[0,1]$, $\Sigma$ is also restricted to take values in $[0,1]$. The (constrained) Hamiltonian ${\cal H}_{\rm A}^{\circ}$ is defined by
${\cal H}_{\rm A}^{\circ}(S)=\dfrac{1}{2}\sup_{u\in[0,1]:u^{2}=S}\big{\{}-u^{-2}\big{\}}=-\frac{1}{2S},\;\text{ for }\;u^{\circ}(S):=\sqrt{S},\;S\in[0,1],$
@@
Comparing (4.1) and (4.2), we clearly need to have equality between the optimal controls to ensure that both problems lead to the same results. This necessitates $u^{\circ}(S)=u^{\star}(\gamma)$, for $(\gamma,S)\in\mathbb{R}\times[0,1]$, and we therefore intuit the following one-to-one correspondence:
$S=\begin{cases}|\gamma|^{-1/2}&\text{if }\gamma<-1,\\ 1&\text{if }\gamma\geq-1.\end{cases}$
@@
$\operatorname*{arg\,max}_{S\in[0,1]}\bigg{\{}{\cal H}^{\circ}_{\rm A}(S)+\dfrac{1}{2}\gamma S\bigg{\}}=\operatorname*{arg\,max}_{S\in[0,1]}\big{\{}-S^{-1}+\gamma S\big{\}}=\begin{cases}|\gamma|^{-1/2}&\text{if }\gamma<-1,\\ 1&\text{if }\gamma\geq-1.\end{cases}$
On the other hand, for $S\in[0,1]$ and $\gamma\in\mathbb{R}$,
${\cal H}_{\rm A}(\gamma)-\dfrac{1}{2}\gamma S=\dfrac{1}{2}\sup_{u\in[0,1]}\big{\{}\gamma u^{2}-u^{-2}\big{\}}-\dfrac{1}{2}\gamma S=\begin{cases}-|\gamma|^{1/2}-\dfrac{1}{2}\gamma S&\text{if }\gamma<-1,\\[5.0pt] \dfrac{1}{2}\big{(}\gamma-1-\gamma S\big{)}&\text{if }\gamma\geq-1.\end{cases}$
implying since $S\in[0,1]$ that
$\operatorname*{arg\,min}_{\gamma\in\mathbb{R}}\bigg{\{}{\cal H}_{\rm A}(\gamma)-\dfrac{1}{2}\gamma S\bigg{\}}=-S^{-2}.$
@@ 4.2 A more intricate example
We now consider a model inspired by the demand-response problem introduced and solved by Aïd, Possamaï, and Touzi (2022) [1]. In this model [^2] We will only briefly describe the model here and refer to [1] for further discussion and motivation regarding the problem under consideration., the agent can impact, through an admissible control $\nu:=(\alpha,\beta)\in{\cal U}$, both the mean and the volatility of his electricity consumption for various specific usages, such as heating, air conditioning, etc. His deviation with respect to his original consumption $x_{0}\in\mathbb{R}_{+}$ is denoted $X$, and satisfies:
@@
As usual, it is assumed that the principal observes the output process $X$ in continuous time, and therefore its quadratic variation. However, contrary to the previous example, these observations are not sufficient to deduce the agent’s effective effort.
@@
for some parameters $\lambda_{k}>0$ and $\mu_{k}>0$, $k\in\{1,\dots,d\}$. [^3] Note that here, for computational convenience and illustration purposes, we slightly simplify the model, in particular the sets $A$ and $B$. Indeed, in the original model, these sets are bounded, which will drastically complicates the agent’s problem in our approach. The principal, who represents an electricity producer/supplier, wishes to incentivise the consumer to decrease the mean and the volatility of his consumption during the contracting period, in order to maximise his expected utility, defined as follows:
@@
If we follow the approach developed in [9], and used in [1], the optimal form of contract is identical to the one introduced for the previous example, namely
@@
for $y_{0}\in\mathbb{R}$, $(Z,\Gamma)\in{\cal V}$, both taking values in $\mathbb{R}$, but the agent’s Hamiltonian is now defined for $(x,z,\gamma)\in\mathbb{R}^{3}$ by
@@
Computing the optimiser of the previous Hamiltonian will determine the optimal response to the contract $\xi$ indexed by $(Z,\Gamma)\in{\cal V}$, i.e. for all $k=1,\dots,d$,
$a^{\star}_{k}(z)=-\mu_{k}z\vee 0,\;\text{ and }\;b^{\star}_{k}(\gamma)=(-\lambda_{k}\gamma)^{-1/2},\;\gamma<0.$
We can compute the dynamics of $X$ and $Y^{y_{0},Z,\Gamma}$ under the optimal effort, and rewrite the more standard stochastic control problem of the principal as follows
@@
with state variables’ dynamics
@@
Proceeding now with our approach, the constraint Hamiltonian is given for $(x,z,S)\in\mathbb{R}^{2}\times\mathbb{R}^{\star}_{+}$ by
@@
One can verify that this coincides with the correspondence given by Lemma 3.11. Indeed, we have
$\operatorname*{arg\,max}_{S>0}\bigg{\{}{\cal H}_{\rm A}^{\circ}(x,z,S)+\dfrac{1}{2}\gamma S\bigg{\}}$ $=\operatorname*{arg\,max}_{S>0}\bigg{\{}-\frac{\widebar{\sigma}^{2}}{S}+\gamma S\bigg{\}}=\dfrac{\widebar{\sigma}}{\sqrt{-\gamma}},\quad\gamma<0,$

### v2

@@ 4 Illustrative Examples
In this section, we first present two illustrative examples, in order to highlight the link between Problem 2 and Problem 4, in particular the one-to-one correspondence between the contract’s sensitivity parameter $\Gamma$ in Problem 2 and the quadratic variation density process $\Sigma$ in Problem 4. Additionally, we present a third example, in which Assumption B is not satisfied, leading to a ‘gap’ between the value of the original problem and its first-best counterpart. This ‘counter-example’ raises additional questions on the optimality of the contracts introduced in [9]. The introduction and resolution of these three examples is rather informal, we refer to the previous sections for a more rigorous formulation and the theoretical results. We highlight that in all three examples, the coefficients are trivially Lipschitz continuous and satisfy the usual linear growth condition, thus Assumption A is always satisfied (see for example Touzi (2013) [32, Theorem 2.2 and 3.1]).
@@ 4.1 A first-best case
We first consider a very simple example, in which the agent only controls the volatility of a one-dimensional output process. More precisely, we let
$X_{t}:=x_{0}+\int_{0}^{t}\nu_{s}\mathrm{d}W_{s},\;t\in[0,T],\;x_{0}\in\mathbb{R},$ (4.1)
where $W$ here is a one-dimensional Brownian Motion, and $\nu$ is the agent’s control, assumed to take values in $(0,1]$. Given a contract $\xi$, the agent wishes to maximise the following expected utility,
@@
We first recall that, since the principal observes $X$ in continuous time, she can also compute the associated quadratic variation process. Here, such process follows the dynamics $\mathrm{d}\langle X\rangle_{t}=\nu_{t}^{2}\mathrm{d}t$, $t\in[0,T]$. In other words, its density process $\Sigma$ satisfies $\Sigma_{t}=\nu_{t}^{2}$, $t\in[0,T]$, and since $\nu$ takes values in $(0,1]$, it is clear that, by observing $\Sigma$ in continuous-time, the principal can deduce the agent’s effort $\nu_{t}=\sqrt{\Sigma_{t}}$, $t\in[0,T]$. This example therefore illustrates a first-best scenario, since from her observation, the principal can actually deduce the agent’s effort.
If one follows the approach in [9], slightly adapted for CARA utility function, the optimal form of contract is
@@
for $y_{0}\in\mathbb{R}$, $(Z,\Gamma)$ an $\mathbb{R}^{2}$-valued process, and where the agent’s Hamiltonian is defined by
${\cal H}_{\rm A}(\gamma):=\dfrac{1}{2}\sup_{u\in(0,1]}\big{\{}\gamma u^{2}-u^{-2}\big{\}},\;\gamma\in\mathbb{R}.$
@@
If we follow our approach, we first need to fix the quadratic variation density process $\Sigma\in{\cal S}$. Note here that, since $\Sigma_{t}=|\nu_{t}|^{2}$, $t\in[0,T]$, and $\nu$ takes values in $(0,1]$, $\Sigma$ is also restricted to take values in $(0,1]$. The (constrained) Hamiltonian ${\cal H}_{\rm A}^{\circ}$ is defined by
${\cal H}_{\rm A}^{\circ}(S)=\dfrac{1}{2}\sup_{u\in(0,1]:u^{2}=S}\big{\{}-u^{-2}\big{\}}=-\frac{1}{2S},\;\text{ for }\;u^{\circ}(S):=\sqrt{S},\;S\in(0,1],$
@@
Comparing (4.2) and (4.3), we clearly need to have equality between the optimal controls to ensure that both problems lead to the same results. This necessitates $u^{\circ}(S)=u^{\star}(\gamma)$, for $(\gamma,S)\in\mathbb{R}\times(0,1]$, and we therefore intuit the following correspondence:
$S^{\star}(\gamma):=\begin{cases}|\gamma|^{-1/2}&\text{if }\gamma<-1,\\ 1&\text{if }\gamma\geq-1.\end{cases}$
@@
$\operatorname*{arg\,max}_{S\in(0,1]}\bigg{\{}{\cal H}^{\circ}_{\rm A}(S)+\dfrac{1}{2}\gamma S\bigg{\}}=\operatorname*{arg\,max}_{S\in(0,1]}\big{\{}-S^{-1}+\gamma S\big{\}}=\begin{cases}|\gamma|^{-1/2}&\text{if }\gamma<-1,\\ 1&\text{if }\gamma\geq-1.\end{cases}$
On the other hand, for $S\in(0,1]$ and $\gamma\in\mathbb{R}$, we compute
${\cal H}_{\rm A}(\gamma)-\dfrac{1}{2}\gamma S=\dfrac{1}{2}\sup_{u\in(0,1]}\big{\{}\gamma u^{2}-u^{-2}\big{\}}-\dfrac{1}{2}\gamma S=\begin{cases}-|\gamma|^{1/2}-\dfrac{1}{2}\gamma S&\text{if }\gamma<-1,\\[5.0pt] \dfrac{1}{2}\big{(}\gamma-1-\gamma S\big{)}&\text{if }\gamma\geq-1.\end{cases}$
Since $S\in(0,1]$, we first show that Assumption B is satisfied. Indeed,
$\min_{\gamma\in\mathbb{R}}\left\{{\cal H}_{\rm A}(\gamma)-\dfrac{1}{2}\gamma S\right\}=-\frac{1}{2S}={\cal H}^{\circ}_{\rm A}(S),\quad S\in(0,1],$
with the minimum achieved for $\gamma^{\star}(S)=S^{-2}$. This is inline with Lemma 3.13.
@@ 4.2 A more intricate example
We now consider a model inspired by the demand-response problem introduced and solved by Aïd, Possamaï, and Touzi (2022) [1]. In this model [^3] We will only briefly describe the model here and refer to [1] for further discussion and motivation., the agent can impact, through an admissible control $\nu:=(\alpha,\beta)\in{\cal U}$, both the mean and the volatility of his electricity consumption for various specific usages, such as heating, air conditioning, etc. His deviation with respect to his original consumption $x_{0}\in\mathbb{R}_{+}$ is denoted $X$, and satisfies:
@@
As usual, it is assumed that the principal observes the output process $X$ in continuous time, and therefore its quadratic variation. However, contrary to the previous example, these observations are not sufficient to deduce the agent’s effective effort, notably because of the dimension of the underlying noise compare to the observed processes.
@@
for some parameters $\lambda_{k}>0$ and $\mu_{k}>0$, $k\in\{1,\dots,d\}$. [^4] Note that here, for computational convenience and illustration purposes, we slightly simplify the model, in particular the sets $A$ and $B$. Indeed, in the original model, these sets are bounded, which will complicate the agent’s problem in our approach. The principal, who represents an electricity producer/supplier, wishes to incentivise the consumer to decrease the mean and the volatility of his consumption during the contracting period, in order to maximise his expected utility, defined as follows:
@@
If we follow the approach developed in [9], and used in [1], the optimal form of contract is given by
@@
for $y_{0}\in\mathbb{R}$, a pair of $\mathbb{R}$-valued processes $(Z,\Gamma)$, and where the agent’s Hamiltonian is defined for $(x,z,\gamma)\in\mathbb{R}^{3}$ by
@@
As usual, computing the optimiser of the Hamiltonian will determine the agent’s best response functions, i.e.
$a^{\star}_{k}(z)=-\mu_{k}z\vee 0,\;\text{ and }\;b^{\star}_{k}(\gamma)=(-\lambda_{k}\gamma)^{-1/2},\;\gamma<0,\quad k=1,\dots,d.$
We can then reformulate the principal’s problem as follows,
@@
with dynamics of state variables $X$ and $Y^{y_{0},Z,\Gamma}$ under the optimal efforts given by
@@
Alternatively, if we proceed with our approach, the constraint Hamiltonian is given for $(x,z,S)\in\mathbb{R}^{2}\times\mathbb{R}^{\star}_{+}$ by
@@
One can verify that this coincides with the correspondence given by Lemmas 3.11 and 3.13. Indeed, we have
$S^{\star}(\gamma):=\operatorname*{arg\,max}_{S>0}\bigg{\{}{\cal H}_{\rm A}^{\circ}(x,z,S)+\dfrac{1}{2}\gamma S\bigg{\}}$ $=\operatorname*{arg\,max}_{S>0}\bigg{\{}-\frac{\widebar{\sigma}^{2}}{S}+\gamma S\bigg{\}}=\dfrac{\widebar{\sigma}}{\sqrt{-\gamma}},\quad\gamma<0,$
@@
In particular, Assumption B is satisfied since
$\min_{\gamma<0}\bigg{\{}{\cal H}_{\rm A}(x,z,\gamma)-\dfrac{1}{2}\gamma S\bigg{\}}=\kappa x+\dfrac{1}{2}\widebar{\mu}(z\wedge 0)^{2}-\dfrac{\widebar{\sigma}^{2}}{S}={\cal H}_{\rm A}^{\circ}(x,z,S),\quad S>0,$
and the minimum is achieved at $\gamma^{\star}(S)$.
4.3 A ‘counter-example’
In this example, we highlight that Assumption B may not always be satisfied, and that, in this case, there may be a difference between:
  • the value of the ‘first-best’ formulation, namely $V^{\circ}_{\rm P}$ in Problem 3, or equivalently $\widetilde{V}^{\circ}_{\rm P}$ in Problem 4,
  • the value of the original problem when restricting to contracts of the form (2.11), i.e. $\widetilde{V}_{\rm P}$ in Problem 2.
For this, we consider the controlled state equation (4.1) driven by a one-dimensional Brownian Motion $W$ as in the first example in Section 4.1, but where now the control process $\nu$ can take values in $U:=[-1,1]$. We note that achievable quadratic variations take values in $[0,1]$. We assume further that the agent’s cost function is given by $c(u)=1-u^{4}$ for $u\in[-1,1]$, thus non-negative and bounded as required.
We first highlight that Assumption B is not satisfied in this framework. Indeed, let $S\in[0,1]$, we have
${\cal H}_{\rm A}^{\circ}(S):=\sup_{u\in U:u^{2}=S}\big{\{}-c(u)\big{\}}=\max\Big{\{}\big{(}\sqrt{S}\big{)}^{4}-1,\big{(}-\sqrt{S}\big{)}^{4}-1\Big{\}}=S^{2}-1,\quad\text{ for }\;u^{\circ}(S)=\pm\sqrt{S}.$ (4.5)
The ‘constraint’ Hamiltonian ${\cal H}_{\rm A}^{\circ}$ is proper, continuous, but convex instead of concave, implying that Assumption B is not satisfied for all $S\in[0,1]$. Indeed, for any $\gamma\in\mathbb{R}$, we have
${\cal H}_{\rm A}(\gamma)$ $:=\sup_{u\in U}\bigg{\{}\dfrac{1}{2}\gamma u^{2}-c(u)\bigg{\}}=\sup_{S\in[0,1]}\bigg{\{}\sup_{u\in[-1,1]:u^{2}=S}\bigg{\{}\dfrac{1}{2}\gamma u^{2}+u^{4}\bigg{\}}\bigg{\}}-1=\sup_{S\in[0,1]}\bigg{\{}\dfrac{1}{2}\gamma S+S^{2}\bigg{\}}-1$
=$=\begin{cases}\gamma/2&\mbox{ if }\gamma>-2,\text{ for }S^{\star}(\gamma)=1,\text{ i.e. }u^{\star}(\gamma)\in\{-1,1\};\\ -1&\mbox{ if }\gamma=-2,\text{ for }S^{\star}(\gamma)\in\{0,1\},\text{ i.e. }u^{\star}(\gamma)\in\{-1,0,1\};\\ -1&\mbox{ if }\gamma<-2,\text{ for }S^{\star}(\gamma)=0,\text{ i.e. }u^{\star}(\gamma)=0.\end{cases}$ (4.6)
Computing the convex conjugate of ${\cal H}_{\rm A}$, we obtain
$\inf_{\gamma\in\mathbb{R}}\bigg{\{}{\cal H}_{\rm A}(\gamma)-\dfrac{1}{2}\gamma S\bigg{\}}=S-1,\quad\text{ for }\gamma^{\star}=-2.$
We thus have
$\inf_{\gamma\in\mathbb{R}}\bigg{\{}{\cal H}_{\rm A}(\gamma)-\dfrac{1}{2}\gamma S\bigg{\}}-{\cal H}_{\rm A}^{\circ}(S)=S-S^{2},$
which is always non-negative, but (strictly) positive on $(0,1)$. In other words, Assumption B is not verified, unless we restrict the study to $S\in\{0,1\}$. Intuitively, if the principal’s optimal choice for the quadratic variation is at either $S=0$ or $S=1$, then no issue should arise, since Assumption B is satisfied at these two points. However, if the optimal value for the quadratic variation lies in the interior of the interval $(0,1)$, the ‘first-best’ value should exceed the value obtained by restricting to contracts of the form (2.11). Indeed, as highlighted above when computing the maximisers of the Hamiltonian ${\cal H}_{\rm A}$, these contracts can only incentivise optimal efforts taking values in $\{-1,0,1\}$, thus leading to a quadratic variation taking values in $\{0,1\}$ only. We confirm this intuition below.
We consider the following criterion for the principal, in which the cost function $c_{\rm P}$ will be chosen later,
$J_{\rm P}(\xi,\Sigma):=\mathbb{E}\bigg{[}X_{T}-\xi-\int_{0}^{T}c_{\rm P}\big{(}\Sigma_{t}\big{)}\mathrm{d}t\bigg{]}.$ (4.7)
In the ‘first-best’ formulation, we know that the optimal contract form is given by
$\xi=y_{0}-\int_{0}^{T}{\cal H}_{\rm A}^{\circ}(\Sigma_{t})\mathrm{d}t+\int_{0}^{T}Z_{t}\mathrm{d}X_{t},$
where the pair of processes $(Z,\Sigma)$ has to be chosen by the principal. The optimal efforts are given by the maximiser of the constrained Hamiltonian ${\cal H}^{\circ}_{\rm A}$, here $\nu_{t}^{\circ}=\pm\sqrt{\Sigma_{t}}$, which leads to ${\cal H}_{\rm A}^{\circ}(\Sigma_{t})=\Sigma_{t}^{2}-1$, $t\in[0,T]$, as shown in (4.5). Using the dynamics of $X$ and $\xi$ under these optimal efforts, we can compute
$J_{\rm P}(\xi,\Sigma)=x_{0}-y_{0}-T+\mathbb{E}\bigg{[}\int_{0}^{T}\big{(}\Sigma_{t}^{2}-c_{\rm P}(\Sigma_{t})\big{)}\mathrm{d}t\bigg{]}.$
By pointwise optimisation, the optimal $\Sigma$ should be given by
$\Sigma_{t}^{\star}\in\operatorname*{arg\,max}_{S\in[0,1]}\big{\{}S^{2}-c_{\rm P}(S)\big{\}},\quad t\in[0,T].$
Now, by choosing the cost function $c_{\rm P}(S)=S^{3}$, we obtain a maximum at $4/27$ achieved for $S^{\star}=2/3\in(0,1)$. In other words, the optimal quadratic variation is neither 00 or $1$, and one can thus expect a duality gap in this case.
Indeed, one can also compute the principal’s value when she can only offer contracts of the form (2.11). In this case, we still consider the criterion (4.7), but with the difference that $\Sigma$ is not directly chosen by the principal. In particular, given a contract of the form (2.11) indexed through a parameter $\Gamma$, we have $\Sigma_{t}=|u^{\star}(\Gamma_{t})|^{2}$, $t\in[0,T]$, where $u^{\star}$ is defined through (4.3) as the maximiser of the Hamiltonian ${\cal H}_{\rm A}$. Using this contract form and the associated optimal efforts, we can then compute
$J_{\rm P}\big{(}\xi,|u^{\star}(\Gamma_{\cdot})|^{2}\big{)}=x_{0}-y_{0}+\mathbb{E}\bigg{[}\int_{0}^{T}\bigg{(}{\cal H}_{\rm A}(\Gamma_{t})-\dfrac{1}{2}\Gamma_{t}|u^{\star}(\Gamma_{t})|^{2}-c_{\rm P}\big{(}|u^{\star}(\Gamma_{t})|^{2}\big{)}\bigg{)}\mathrm{d}t\bigg{]},$
where now the process $\Gamma$ has to be chosen optimally. Again by pointwise optimisation, we should have
$\Gamma_{t}^{\star}\in\operatorname*{arg\,max}_{\gamma\in\mathbb{R}}\bigg{\{}{\cal H}_{\rm A}(\gamma)-\dfrac{1}{2}\gamma|u^{\star}(\gamma)|^{2}-c_{\rm P}\big{(}|u^{\star}(\gamma)|^{2}\big{)}\bigg{\}},\quad t\in[0,T].$
By choosing $c_{\rm P}(S)=S^{3}$ as before, the maximisation problem becomes
$\max_{\gamma\in\mathbb{R}}\bigg{\{}{\cal H}_{\rm A}(\gamma)-\dfrac{1}{2}\gamma|u^{\star}(\gamma)|^{2}-|u^{\star}(\gamma)|^{6}\bigg{\}}=-1,\quad\text{for }\;\gamma^{\star}=-2.$
Therefore, while the value in the ‘first-best’ case is given by $x_{0}-y_{0}-23T/27$, the maximum possible value achievable with the contracts (2.11) in the original ‘second-best’ problem is lower, equal to $x_{0}-y_{0}-T$.
This illustrative example highlights that if Assumption B does not hold, the contract form (2.11) may fail to achieve the ‘first-best’ value. As mentioned in Remark 3.12, without Assumption B, the proof of optimality of the contract form (2.11) via 2BSDEs may also break down. Therefore, while this example illustrates a difference between the value of the original problem when restricting attention to contracts of the form (2.11) (namely, Problem 2) and the value of the ‘first-best’ reformulation (Problems 3 and 4), it remains unclear whether this reflects an intrinsic ‘gap’ between Problem 1 and Problem 3, or whether it is possible to design a new class of penalisation contracts to bridge this gap. We believe this is a very interesting question, which we leave for further research.